Consider the following three functions:

(A) "x^2"

(B) "x\cos{x}"

(C) "e^{x}"

and select the correct option for the following:


  1. The function (A, B or C) is an even function
  2. (A, B or C) is an odd function while the function
  3. (A, B or C) is neither odd nor even.
1
Expert's answer
2021-09-21T03:06:05-0400

A) "f(x) = x^2" "f(-x) = (-x)^2=x^2" -> this is even function

B) "f(x) = xcosx" "f(-x) = (-x)cos(-x) = -xcosx = -f(x)" -> this is odd function

C) "f(x) = e^x" "f(-x) = e^{-x} = \large\frac{1}{e^x}=\frac{1}{f(x)}" -> this function is neither odd nor even.


  1. The function (A, B or C) is an even function (true) A is even function
  2. (A, B or C) is an odd function (true) B is odd function
  3. (A, B or C) is neither odd nor even. (true) C is neither odd nor even.